the sum of three consecutive even intergers is fourteen less than twice the smallest of these intergers
Answer by checkley75(3666)

X+(X+2)+(X+4)=2X+14
3X+6=2X+14
3X-2X=14-6
X=8 ANSWER FOR THE SMALLEST NUMBER.
PROOF
8+10+12=2*8+14
30=16+14
30=30